The horizontal and vertical displacement a of a projectile are given by x=3tmetees and y=4t-5t^2metre. The initial velocity of projectile is options: 3 4 5 6

Grade:12

1 Answers

Arun
6 years ago
dx/dt = 3
dy/dt = 4 - 10t
In starting, t= 0
Hence dy/dt = 4
Hence initial velocity = sqrt(3²+4²) = sqrt(25) = 5 m/sec
 
Regards
Arun (askIITians forum expert)